3. What volume of 40.0% NaNO3 solution contains 0.15 mole of NaNO3? Density 1.32 g/mL
Answer #1

40.0% NaNO3 means

40.0 grams of NaNO3 present in 100 mL of solution

Density = 1.32 g/mL

i.e mass of 1 mL of NaNO3 = 1.32 grams

mass of 100 mL of NaNO3 = 1.32x100 = 132.0 grams

40% of 132 grams = 52.8 grams of NaNO3

number of moles of NaNO3= 0.15 moles

molar mass of NaNO3= 85.0 g/mole

mass of 0.15 moles of NaNO3 = 0.15x85.0=12.75 grams of NaNO3

52.8/132 = 12.75/x grams solution

x= 31.875 grams of solution

mass of solution = 31.87 grams

volume of 1.32 grams of solution = 1 mL

volume of 31.87 grmas of solution = ?

                                                    = 31.87x1/1.32 = 24.14 mL

volume of solution = 24.14 mL